Grade 12th passPhysical Chemistry

What is no. Of moles of O-atom in 126 amu of HNO3?.....

rana
9 Years agoGrade 12th pass

3 Answers

Siddhant mudholkar
9 Years ago
Molecular mass of HNO3=63NO. OF MOLES OF HNO3 IN 126u=126÷63=2molesNo.of moles of oxygen in 1 mole of HNO3=3Therefore answer is 6 moles of oxygen
Sheetal kumari
9 Years ago
Thanks for askingTo find the number of moles we divide given mass of a substance by its atomic mass.Thus we have given 126 amu of HNO3 we can find its number of moles by dividing 126 by 63 (which is the moleculr mass of HNO3)=2 molesThus if 1 mole of HNO3 contains 3 moles of oxygen then its 2 moles contains 6 moles of oxygenPlease kindly approved me
Lavish Sachdeva
7 Years ago
126 amu HNO3 (Molecular mass is 63 amu) 
Means there Are 2molecules of it
Hence 6 molecules of O=6/NA moles of O.